OpenStudy (cutepuppies):

Which statements are true about the place values in the number 99.999? Choose all answers that are correct. A. The 9 in the hundredths place is 1/10 the value of the 9 in the tenths place. B. The 9 in the tenths place is 10 times the value of the 9 in the hundredths place. C. The 9 in the hundredths place is 10 times the value of the 9 in the tenths place. D. The 9 in the tenths place is 1/10 the value of the 9 in the hundredths place.

OpenStudy (cutepuppies):

@pooja195

OpenStudy (cutepuppies):

@pooja195

OpenStudy (cutepuppies):

@pooja195

OpenStudy (cutepuppies):

@triciaal

jhonyy9 (jhonyy9):

my opinion A. and B. are correct A. because 0,09=0,9 * 1/10 = 0,9*0,1=0,09 and for B. 0,9 = 0,09*10=0,9 - hope these will help you

OpenStudy (triciaal):